OpenStudy (anonymous):

Consider the functions f(x) = 2(x − 9) and g(x) = 4x. What is the value of (f − g)(5)? 12 28 20 −28

OpenStudy (zehanz):

Just remember: (f-g)(5)=f(5)-g(5)

OpenStudy (anonymous):

So what do i plug in?

OpenStudy (zehanz):

Plug in 5 in both of the functions and subtract the outcomes.

OpenStudy (anonymous):

so, f(5) = 2(5-9) and g(5)=4 times 5?

OpenStudy (anonymous):

now i am confused

OpenStudy (anonymous):

simplify f(5) = ??? g(5) = ???

OpenStudy (anonymous):

you got it right earlier... just need to simplify....

OpenStudy (anonymous):

ohhh...

OpenStudy (zehanz):

@Andrewisbadatmath: right, and if you've found f(5) and g(5), you calculate their difference: f(5) - g(5)

OpenStudy (anonymous):

ok thank you

OpenStudy (zehanz):

YW!

OpenStudy (anonymous):

so it is -12?

OpenStudy (anonymous):

no i meant 12

OpenStudy (zehanz):

f(5) = 2(5-9) = 2*(-4)=... g(5) = 4*5 = ... f(5) - g(5) = ... (not 12)

OpenStudy (zehanz):

You will get -8 - 20 = ...